# Neural Dynamic Data Valuation: A Stochastic Optimal Control Approach

### v4

@@ 1 Introduction
Data has become a central asset in the modern economy, functioning as intellectual property that drives value creation in data marketplaces [1, 2, 3]. The intrinsic worth of data facilitates sharing, exchange, and reuse across organizations, businesses, and research communities. This value arises from multiple intertwined factors such as data quality, market demand, computational sophistication of downstream tasks, and data integrity, all of which determine the reliability and utility of data across domains.
@@
Earlier research on data valuation has evolved from model-specific influence estimation to cooperative game formulations and recent efficiency-driven approximations. The influence-based approach [4] estimates each sample’s importance through sensitivity analysis of model parameters but struggles to capture complex dependencies among data points. The game-theoretic framework [5, 6, 7, 8, 9] assigns value through average marginal contributions, offering strong fairness principles but requiring exponentially many evaluations. To mitigate this burden, efficient approximation techniques have been introduced, such as neighbor-based models [10, 11, 12] and ensemble reweighting strategies [13, 14], yet these often compromise fairness or interpretability. Further, explainability-oriented studies [15] enhance transparency in local model reasoning but do not extend to global data valuation. Despite these advances, existing methods remain limited by three unresolved issues: excessive computational cost that restricts scalability, biased or inconsistent value allocation across heterogeneous data, and lack of transparency regarding how value evolves during learning. These unresolved problems motivate a unified framework that models data value dynamically, enabling efficient, fair, and interpretable estimation at scale [1, 2, 3].
@@ 2.1 Dynamics and Optimal Control Theory
The dynamical interpretation of deep neural networks (DNNs) as continuous-time systems has introduced new theoretical perspectives on learning and optimization. Foundational studies view deep residual architectures as discrete approximations of ordinary differential equations (ODEs), bridging neural network propagation and numerical integration theory [16, 17, 18]. This connection enables analysis in the continuum limit through Wasserstein geometry, which links gradient flow and optimal transport [19]. The resulting framework naturally extends to mean-field control formulations for learning dynamics [20, 21].
Building on this perspective, several optimization algorithms have been reformulated as control problems [22, 23]. Stochastic gradient descent (SGD) has been analyzed through the lens of stochastic dynamical systems, where Langevin and Lévy dynamics describe implicit regularization effects [24, 25]. Further, studies of Gram matrix evolution have demonstrated convergence and global optimality properties [26, 27], providing guidance for adaptive tuning of hyperparameters such as learning rate and batch size [28, 29]. Collectively, these results suggest that the theory of dynamic systems and optimal control offers a principled foundation for reinterpreting neural learning and motivates the control-based formulation adopted in this work.
@@ 2.2 Data Valuation
Data valuation has emerged as a key research topic to quantify each sample’s contribution to model performance. Early approaches such as Leave-One-Out (LOO) [4] and Data Shapley [10, 5] estimate marginal contributions by iteratively removing data points and observing model performance changes. Although theoretically grounded, these methods suffer from exponential computational cost. Extensions like Beta Shapley [9] and Data Banzhaf [30] improve fairness modeling but remain computationally demanding. Influence-based estimators [31] leverage gradient approximations to reduce complexity, yet often lose robustness in nonconvex learning settings.
To enhance scalability, proxy and ensemble-based techniques were introduced. KNNShapley [11] employs a $k$-Nearest Neighbor classifier as a learning-agnostic surrogate, while AME [13] and Data-OOB [14] adopt ensemble reweighting and out-of-bag estimators for efficiency. These methods, however, still struggle to maintain fairness and transparency. Algorithm-agnostic formulations, such as volume-based valuation [32] and Wasserstein-distance metrics [33], bypass model retraining but lack the ability to detect label errors or explain valuation outcomes.
Beyond standard utility estimation, marginal contribution analysis has been extended to feature attribution [34, 35], interpretability [36, 37], and collaborative learning [38, 39]. Further refinements [40, 41] relax classical Shapley axioms to improve efficiency, while reinforcement learning-based estimators [42] model data usage likelihood during training. Despite these developments, current frameworks still face three inherent constraints: high computational cost, unfair valuation under data heterogeneity, and limited interpretability.
@@ 3 Preliminaries and Problem formulation
which quantifies the change in utility when the $i$-th point is excluded from the full dataset. This metric approximates the influence of a data point but can yield inaccurate values close to zero in practice [43].
@@ 4 Method
The proposed Neural Dynamic Data Valuation (NDDV) framework reformulates the data valuation problem as a stochastic optimal control process, where data contribution evolves continuously within the learning dynamics of a neural network. NDDV comprises three tightly integrated modules that operate coherently to achieve these objectives. The first module, the Dynamic State Encoder, transforms data samples into a latent representation governed by stochastic differential equations, which model their dynamic interactions with the evolving learning system [24, 23]. The second module, the Fairness-Aware Mean-Field Controller, defines a mean-field potential that regulates global fairness across samples by adjusting their control weights according to their statistical influence, ensuring equitable value allocation even under data heterogeneity [44, 45]. The third module, the Interpretable Neural Valuator, leverages KANs with Matérn kernels [46, 47] to map the evolving latent states into valuation scores, preserving the underlying dynamics while offering transparency in the contribution estimation. Collectively, these modules constitute a unified control-driven valuation framework that jointly ensures efficiency, fairness and interpretability. This design provides a principled and scalable solution to overcome the core deficiencies of prior valuation paradigms and enables reliable deployment of data valuation in large-scale machine learning systems.
@@ 4.1 Learning Dynamic Valuation Trajectories via Stochastic Control
To address the inefficiency and static limitations of conventional data valuation, we formulate valuation as a stochastic optimal control problem [48, 49]. This dynamic view models how data points jointly influence learning over time, where each point follows a stochastic trajectory capturing its evolving contribution (see Fig. 2).
@@
Applying the Stochastic Maximum Principle (SMP) [50], we introduce the Hamiltonian:
@@
The coupled FSDE–BSDE system forms a McKean–Vlasov FBSDE [51, 44, 52], and the optimal control $\psi_{t}^{*}$ satisfies:
@@
linking valuation to the gradient of the cost function [53, 54].
@@
This stochastic dynamic formulation eliminates combinatorial retraining, models temporal evolution of data influence, and offers interpretability across layers and epochs, the capabilities absent in static valuation heuristics [55, 56].
@@ 4.2 Fairness-Aware Valuation via Reweighting and Mean-Field Dynamics
Adaptive Reweighting via Meta-Learned Weights. Each data point is dynamically reweighted by a meta-learned function $V(\Phi_{i}(X_{T});\theta)$ inspired by Stackelberg game theory [45]. The control objective becomes:
@@
solved efficiently through a single-loop online update [57, 58]. This procedure amplifies representative and fair samples while reducing bias from noisy or minority points.
Mean-Field Interaction and Weighted Aggregation. Group fairness is enforced by embedding reweighted data into a mean-field dynamic system [44]:
@@
With a linear–quadratic drift [59, 45, 60],
@@
Formal Fairness Guarantee. Following [5, 61], NDDV satisfies symmetry, monotonicity, and null-player properties. Fairness is quantified by the Equal Opportunity (EOp) and Equalized Odds (EOdds) scores [62]:
@@ 4.3 Learning Interpretable Utility Function
Inspired by KANs [46], we replace dense weight matrices with univariate basis functions, expressing $\psi$ and $\mathcal{V}$ as compositions of nonlinear mappings for improved interpretability and analytical control. The control strategy $\psi$ is represented as:
@@
Because Gaussian RBFs can over-smooth noisy distributions, we adopt a **Matérn kernel** [47] to improve local adaptivity:
@@ 6.1 Experimental Setup
Datasets. We evaluate our approach using six publicly available datasets provided in OpenDataVal [63], several of which have been widely adopted in prior work [5, 9]. These datasets span diverse domains—including image, text, and tabular data with varying degrees of class imbalance, feature sparsity, and task complexity, as detailed in Table I. This diversity ensures a comprehensive assessment of both the scalability and generalizability of our method. Moreover, the inclusion of datasets previously used in influential studies facilitates direct comparisons and fair benchmarking.
Baselines. We compare NDDV against eight representative data valuation methods: LOO [4], DataShapley [5], BetaShapley [9], DataBanzhaf [30], InfluenceFunction [31], KNNShapley [11], AME [13], and Data-OOB [14]. These methods cover a wide spectrum of valuation paradigms, including exact and approximate game-theoretic methods, influence-based techniques, and ensemble-based approaches. This selection ensures a balanced evaluation across accuracy, efficiency, and robustness. For fairness, all baselines are run with the same or greater number of utility evaluations than NDDV.
@@
2dplanes [64] 40768 10 2 0.499 Tabular
electricity [65] 38474 6 2 0.5 Tabular
fired [64] 40768 10 2 0.498 Tabular
BBC [66] 2225 768 5 0.17 Text
IMDB [67] 50000 768 2 0.5 Text
STL10 [68] 5000 96 10 0.01 Image
CIFAR10 [69] 50000 2048 10 0.1 Image
@@ 6.4 Robustness Against Data Corruption and Manipulation
High-Value Data Removal. Following [5, 14], we iteratively remove the most valuable samples and re-train the model (Fig. 7, second column). NDDV causes the steepest test accuracy decline, confirming its precise identification of influential data, while AME and KNNShapley show weaker prioritization.
@@ 6.5 Ablation Study
When the mean-field controller was disabled, both Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD) increased significantly, showing that fairness improvement in NDDV arises directly from the adaptive mean-field interactions. Removing the dynamic state encoder led to a noticeable drop in F1-score and rank stability, as the model could no longer capture temporal dependencies between evolving data representations. Similarly, substituting the interpretable neural valuator with a standard multilayer perceptron degraded Rank Consistency (RC) and Spearman Correlation (SC) by over 20%, validating that the Kolmogorov–Arnold representation [46, 47] enhances the consistency and interpretability of valuation.
@@ 6.6 Fairness Evaluation
This experiment investigates how effectively NDDV mitigates valuation bias and ensures equitable contribution assessment across heterogeneous data groups. The study analyzes feature bias and label noise bias following the established evaluation design in fairness-aware machine learning [62, 70, 71, 72]. The quantitative results are visualized in Figure 9.
In the feature-bias test, the data distribution was intentionally skewed by introducing correlations between sensitive attributes (e.g., gender or age) and predictive features. Under this setting, traditional valuation approaches—such as InfluenceFunction [31] and DataShapley [5]—exhibited amplified bias, reflected by high Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD) values. In contrast, NDDV maintained near-zero parity gaps due to its mean-field controller, which regularizes global data interactions through adaptive reweighting.
In the label-noise scenario, 15% of the labels were randomly flipped to simulate annotation errors. Here, Data-OOB [14] and AME [13] partially reduced bias but suffered from unstable group-level valuations. NDDV achieved the lowest SPD and EOD across all noise levels, demonstrating its robustness to both systematic and random unfairness sources. The fairness improvement arises from the dynamic correction mechanism, which adjusts each sample’s influence as the control process evolves, ensuring balanced value propagation even when the dataset becomes corrupted.

### v5

@@ 1 Introduction
Data has become a central asset in the modern economy, functioning as intellectual property that drives value creation in data marketplaces [agarwal2019, tian2022private, pei2020survey]. The intrinsic worth of data facilitates sharing, exchange, and reuse across organizations, businesses, and research communities. This value arises from multiple intertwined factors such as data quality, market demand, computational sophistication of downstream tasks, and data integrity, all of which determine the reliability and utility of data across domains.
@@
Earlier research on data valuation has evolved from model-specific influence estimation to cooperative game formulations and recent efficiency-driven approximations. The influence-based approach [koh2017understanding] estimates each sample’s importance through sensitivity analysis of model parameters but struggles to capture complex dependencies among data points. The game-theoretic framework [ghorbani2019, shapley1953value, ghor2020dis, schoch2022cs, kwon2022beta] assigns value through average marginal contributions, offering strong fairness principles but requiring exponentially many evaluations. To mitigate this burden, efficient approximation techniques have been introduced, such as neighbor-based models [jia2019, jia2019b, kwon2021efficient] and ensemble reweighting strategies [lin2022measuring, kwon2023dataoob], yet these often compromise fairness or interpretability. Further, explainability-oriented studies [chen2022explaining] enhance transparency in local model reasoning but do not extend to global data valuation. Despite these advances, existing methods remain limited by three unresolved issues: excessive computational cost that restricts scalability, biased or inconsistent value allocation across heterogeneous data, and lack of transparency regarding how value evolves during learning. These unresolved problems motivate a unified framework that models data value dynamically, enabling efficient, fair, and interpretable estimation at scale [agarwal2019, tian2022private, pei2020survey].
@@ 2.1 Dynamics and Optimal Control Theory
The dynamical interpretation of deep neural networks (DNNs) as continuous-time systems has introduced new theoretical perspectives on learning and optimization. Foundational studies view deep residual architectures as discrete approximations of ordinary differential equations (ODEs), bridging neural network propagation and numerical integration theory [weinan2017proposal, he2016deep, lu2017beyond]. This connection enables analysis in the continuum limit through Wasserstein geometry, which links gradient flow and optimal transport [JMLR:v20:16-243]. The resulting framework naturally extends to mean-field control formulations for learning dynamics [hu2017control, han2018mean].
Building on this perspective, several optimization algorithms have been reformulated as control problems [li2017maximum, li2018optimal]. Stochastic gradient descent (SGD) has been analyzed through the lens of stochastic dynamical systems, where Langevin and Lévy dynamics describe implicit regularization effects [pavliotis2014stochastic, simsekli2019tail]. Further, studies of Gram matrix evolution have demonstrated convergence and global optimality properties [du2018gradienta, du2018gradientb], providing guidance for adaptive tuning of hyperparameters such as learning rate and batch size [li2017stochastic, an2018stochastic]. Collectively, these results suggest that the theory of dynamic systems and optimal control offers a principled foundation for reinterpreting neural learning and motivates the control-based formulation adopted in this work.
@@ 2.2 Data Valuation
Data valuation has emerged as a key research topic to quantify each sample’s contribution to model performance. Early approaches such as Leave-One-Out (LOO) [koh2017understanding] and Data Shapley [jia2019, ghorbani2019] estimate marginal contributions by iteratively removing data points and observing model performance changes. Although theoretically grounded, these methods suffer from exponential computational cost. Extensions like Beta Shapley [kwon2022beta] and Data Banzhaf [wang2022data] improve fairness modeling but remain computationally demanding. Influence-based estimators [feldman2020neural] leverage gradient approximations to reduce complexity, yet often lose robustness in nonconvex learning settings.
To enhance scalability, proxy and ensemble-based techniques were introduced. KNNShapley [jia2019b] employs a $k$-Nearest Neighbor classifier as a learning-agnostic surrogate, while AME [lin2022measuring] and Data-OOB [kwon2023dataoob] adopt ensemble reweighting and out-of-bag estimators for efficiency. These methods, however, still struggle to maintain fairness and transparency. Algorithm-agnostic formulations, such as volume-based valuation [xu2021validation] and Wasserstein-distance metrics [just2023lava], bypass model retraining but lack the ability to detect label errors or explain valuation outcomes.
Beyond standard utility estimation, marginal contribution analysis has been extended to feature attribution [lundberg2017unified, covert2021explaining], interpretability [stier2018analysing, ghorbani2020neuron], and collaborative learning [sim2020collaborative, xu2021gradient]. Further refinements [yan2021if, rozemberczki2022shapley] relax classical Shapley axioms to improve efficiency, while reinforcement learning-based estimators [yoon2020data] model data usage likelihood during training. Despite these developments, current frameworks still face three inherent constraints: high computational cost, unfair valuation under data heterogeneity, and limited interpretability.
@@ 3 Preliminaries and Problem formulation
which quantifies the change in utility when the $i$-th point is excluded from the full dataset. This metric approximates the influence of a data point but can yield inaccurate values close to zero in practice [basu2020influence].
@@ 4 Method
The proposed Neural Dynamic Data Valuation (NDDV) framework reformulates the data valuation problem as a stochastic optimal control process, where data contribution evolves continuously within the learning dynamics of a neural network. NDDV comprises three tightly integrated modules that operate coherently to achieve these objectives. The first module, the Dynamic State Encoder, transforms data samples into a latent representation governed by stochastic differential equations, which model their dynamic interactions with the evolving learning system [pavliotis2014stochastic, li2018optimal]. The second module, the Fairness-Aware Mean-Field Controller, defines a mean-field potential that regulates global fairness across samples by adjusting their control weights according to their statistical influence, ensuring equitable value allocation even under data heterogeneity [carmona2013control, bensoussan2017linear]. The third module, the Interpretable Neural Valuator, leverages KANs with Matérn kernels [liu2024kan, rasmussen2003gaussian] to map the evolving latent states into valuation scores, preserving the underlying dynamics while offering transparency in the contribution estimation. Collectively, these modules constitute a unified control-driven valuation framework that jointly ensures efficiency, fairness and interpretability. This design provides a principled and scalable solution to overcome the core deficiencies of prior valuation paradigms and enables reliable deployment of data valuation in large-scale machine learning systems.
@@ 4.1 Learning Dynamic Valuation Trajectories via Stochastic Control
To address the inefficiency and static limitations of conventional data valuation, we formulate valuation as a stochastic optimal control problem [casert2024learning, gao2024learning]. This dynamic view models how data points jointly influence learning over time, where each point follows a stochastic trajectory capturing its evolving contribution (see Fig. 2).
@@
Applying the Stochastic Maximum Principle (SMP) [kushner1965stochastic], we introduce the Hamiltonian:
@@
The coupled FSDE–BSDE system forms a McKean–Vlasov FBSDE [carmona2015forward, carmona2013control, sen2016nonlinear], and the optimal control $\psi_{t}^{*}$ satisfies:
@@
linking valuation to the gradient of the cost function [shrikumar2017learning, ancona2019explaining].
@@
This stochastic dynamic formulation eliminates combinatorial retraining, models temporal evolution of data influence, and offers interpretability across layers and epochs, the capabilities absent in static valuation heuristics [serban_cvodes_2005, jorgensen_2007].
@@ 4.2 Fairness-Aware Valuation via Reweighting and Mean-Field Dynamics
Adaptive Reweighting via Meta-Learned Weights. Each data point is dynamically reweighted by a meta-learned function $V(\Phi_{i}(X_{T});\theta)$ inspired by Stackelberg game theory [bensoussan2017linear]. The control objective becomes:
@@
solved efficiently through a single-loop online update [shu2019meta, yang2024curriculum]. This procedure amplifies representative and fair samples while reducing bias from noisy or minority points.
Mean-Field Interaction and Weighted Aggregation. Group fairness is enforced by embedding reweighted data into a mean-field dynamic system [carmona2013control]:
@@
With a linear–quadratic drift [yong2013linear, bensoussan2017linear, carmona2015mean],
@@
Formal Fairness Guarantee. Following [ghorbani2019, sim2022data], NDDV satisfies symmetry, monotonicity, and null-player properties. Fairness is quantified by the Equal Opportunity (EOp) and Equalized Odds (EOdds) scores [hardt2016equality]:
@@ 4.3 Learning Interpretable Utility Function
Inspired by KANs [liu2024kan], we replace dense weight matrices with univariate basis functions, expressing $\psi$ and $\mathcal{V}$ as compositions of nonlinear mappings for improved interpretability and analytical control. The control strategy $\psi$ is represented as:
@@
Because Gaussian RBFs can over-smooth noisy distributions, we adopt a **Matérn kernel** [rasmussen2003gaussian] to improve local adaptivity:
@@ 6.1 Experimental Setup
Datasets. We evaluate our approach using six publicly available datasets provided in OpenDataVal [jiang2023opendataval], several of which have been widely adopted in prior work [ghorbani2019, kwon2022beta]. These datasets span diverse domains—including image, text, and tabular data with varying degrees of class imbalance, feature sparsity, and task complexity, as detailed in Table I. This diversity ensures a comprehensive assessment of both the scalability and generalizability of our method. Moreover, the inclusion of datasets previously used in influential studies facilitates direct comparisons and fair benchmarking.
Baselines. We compare NDDV against eight representative data valuation methods: LOO [koh2017understanding], DataShapley [ghorbani2019], BetaShapley [kwon2022beta], DataBanzhaf [wang2022data], InfluenceFunction [feldman2020neural], KNNShapley [jia2019b], AME [lin2022measuring], and Data-OOB [kwon2023dataoob]. These methods cover a wide spectrum of valuation paradigms, including exact and approximate game-theoretic methods, influence-based techniques, and ensemble-based approaches. This selection ensures a balanced evaluation across accuracy, efficiency, and robustness. For fairness, all baselines are run with the same or greater number of utility evaluations than NDDV.
@@
2dplanes [feurer-arxiv19a] 40768 10 2 0.499 Tabular
electricity [gama2004learning] 38474 6 2 0.5 Tabular
fired [feurer-arxiv19a] 40768 10 2 0.498 Tabular
BBC [greene2006practical] 2225 768 5 0.17 Text
IMDB [maas2011learning] 50000 768 2 0.5 Text
STL10 [coates2011analysis] 5000 96 10 0.01 Image
CIFAR10 [krizhevsky2009learning] 50000 2048 10 0.1 Image
@@ 6.4 Robustness Against Data Corruption and Manipulation
High-Value Data Removal. Following [ghorbani2019, kwon2023dataoob], we iteratively remove the most valuable samples and re-train the model (Fig. 7, second column). NDDV causes the steepest test accuracy decline, confirming its precise identification of influential data, while AME and KNNShapley show weaker prioritization.
@@ 6.5 Ablation Study
When the mean-field controller was disabled, both Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD) increased significantly, showing that fairness improvement in NDDV arises directly from the adaptive mean-field interactions. Removing the dynamic state encoder led to a noticeable drop in F1-score and rank stability, as the model could no longer capture temporal dependencies between evolving data representations. Similarly, substituting the interpretable neural valuator with a standard multilayer perceptron degraded Rank Consistency (RC) and Spearman Correlation (SC) by over 20%, validating that the Kolmogorov–Arnold representation [liu2024kan, rasmussen2003gaussian] enhances the consistency and interpretability of valuation.
@@ 6.6 Fairness Evaluation
This experiment investigates how effectively NDDV mitigates valuation bias and ensures equitable contribution assessment across heterogeneous data groups. The study analyzes feature bias and label noise bias following the established evaluation design in fairness-aware machine learning [hardt2016equality, chouldechova2017fair, zafar2017fairness, berk2021fairness]. The quantitative results are visualized in Figure 9.
In the feature-bias test, the data distribution was intentionally skewed by introducing correlations between sensitive attributes (e.g., gender or age) and predictive features. Under this setting, traditional valuation approaches—such as InfluenceFunction [feldman2020neural] and DataShapley [ghorbani2019]—exhibited amplified bias, reflected by high Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD) values. In contrast, NDDV maintained near-zero parity gaps due to its mean-field controller, which regularizes global data interactions through adaptive reweighting.
In the label-noise scenario, 15% of the labels were randomly flipped to simulate annotation errors. Here, Data-OOB [kwon2023dataoob] and AME [lin2022measuring] partially reduced bias but suffered from unstable group-level valuations. NDDV achieved the lowest SPD and EOD across all noise levels, demonstrating its robustness to both systematic and random unfairness sources. The fairness improvement arises from the dynamic correction mechanism, which adjusts each sample’s influence as the control process evolves, ensuring balanced value propagation even when the dataset becomes corrupted.
@@ 7 Conclusion
[Uncaptioned image] Zhangyong Liang is working towards the PhD degree at the National Center for Applied Mathematics, Tianjin University. His research interests include AI for science, physics-informed machine learning, neural PDE solvers, data integration, and explainability.
[Uncaptioned image] Ji Zhang is currently a full Professor in Computer Science at the University of Southern Queensland (UniSQ), Australia. He is an RSA Life Fellow, IET Fellow, BCS Fellow, IAAST Fellow, IEEE Senior Member, Australian Endeavour Fellow, Queensland International Fellow (Australia), and Izaak Walton Killam Scholar (Canada). His research interests span big data analytics, data intelligence, intelligent computing, computational intelligence and data privacy. He has published over 380 papers in major peer-reviewed international journals and conferences including TPAMI, TKDE, TKDD, TDSC, TIFS, TIST, CSUR, AAAI, IJCAI, VLDB, CIKM, SIGKDD, ICDE, ICDM, CVPR and WWW.
[Uncaptioned image] Xin Wang received the Ph.D. degree from the University of Edinburgh in 2013. His research mainly focuses on Big Data, Knowledge Engineering and Natural Language Processing. He is a professor in the School of Computer Science and software engineering, Southwest Petroleum University, and serves as a Sichuan Provincial Distinguished Expert, China. He has published papers in the prestigious conferences and journals, including SIGMOD, VLDB, ICDE, CVPR and the IEEE Transactions on Knowledge and Data Engineering, Expert Systems with Applications, and received ICDE 2014 Best Paper Runner-up Award, WISE 2019 Best Paper Runner-up Award, etc.
[Uncaptioned image] Pengfei Zhang received a Ph.D from Beijing University of Posts and Telecommunications in China, with a major in Computer Science. He has served as the reviewer, meta-reviewer or area chair for many top-tier conferences and journals. He has published more than 30 papers in NeurIPS, TKDE, IJCAI, TIFS, AAAI, CIKM and so on. He is also a guest editor for several journals. His current research interest focuses on big data, privacy protection and mobile crowdsensing.
[Uncaptioned image] Zhao Li received his PhD degree in Computer Science from the University of Vermont in 2012, receiving the Excellent Graduate Award. He is currently an Adjunct Professor at Zhejiang University and co-founder of Hangzhou Yugu Technology Co., Ltd. He has published numerous papers in prestigious conferences and journals, including WWW, AAAI, IAAI, ICDE, VLDB, and TKDE. His research interests include machine learning, big data-driven applications, and large-scale graph computing.